OpenStudy (anonymous):

solve for x. (4/6x+3)=(5/5x-8) +1

OpenStudy (anonymous):

\[\frac{ 4 }{ 6x+3 }=\frac{ 5 }{ 5x-8 } +\frac{ 5x-8}{ 5x-8 }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 6x+3 }=\frac{ 5x-3 }{ 5x-8 }\]

OpenStudy (anonymous):

are the steps making sense so far?

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

\[4(5x-8)=(6x+3)(5x-3)\]

OpenStudy (anonymous):

\[20x-32=30x^2-3x-9\]

OpenStudy (anonymous):

\[0=30x^2-23x+23\]

OpenStudy (anonymous):

Next you use quadratic formula. \[x=\frac{ -(-23) +or-\sqrt{(-23)^2-4(30)(23)} }{ 2(30)}\]

OpenStudy (anonymous):

I really appreciate you taking your time to walk me though the process. thank you so much :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

\[\frac{ 23+ or -\sqrt{-2231} }{ 60 }\]

OpenStudy (anonymous):

so the two possible x values are imaginary. \[\frac{ 23 }{ 60 } + \frac{ i \sqrt{2231} }{ 60 } and \frac{ 23}{ 60 }-\frac{ i \sqrt{2231} }{ 60 }\]

OpenStudy (anonymous):

does the imaginary have a value?

OpenStudy (anonymous):

Imaginary numbers do not have values, because they aren't real.

OpenStudy (anonymous):

so if i were to check the answer, it wouldn't really check so it would be no solution?

OpenStudy (anonymous):

There's a solution, but it's imaginary, so you can't really work with it.

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

No problem, glad to help.